OpenStudy (fibonaccichick666):

Hey, can y'all check this: \[\lim_{ {x->-2^{-}}} \frac{|x^2-x-6|}{(x^2-4x+4)}\] The book is saying the answer is \(+\infty \), but I am getting 0.

OpenStudy (anonymous):

Did you factor them both or what?

OpenStudy (fibonaccichick666):

I found for which x the abs val is positive, -2 is the branch pt, then because the fn exists at x=-2 I plugged in the corresponding abs val. The issue is that it is a cusp, but I just can't remember how to get infinity.

OpenStudy (fibonaccichick666):

and, I just did it with factoring and I'm still getting 0.

OpenStudy (fibonaccichick666):

not sure if I'm just stupid, or the book is wrong

OpenStudy (anonymous):

\[ \lim_{ {x->2^{-}}} \frac{|x^2-x-6|}{(x^2-4x+4)}= \frac{|2^2-2-6|}{(2^2-4(2)+4)}=4/0=\infty \]

OpenStudy (fibonaccichick666):

@eliassaab Sorry, missed the -. It's -2

OpenStudy (anonymous):

It is \(+\infty\) since the numerator and the denominator are positive near 2

OpenStudy (nincompoop):

what the f#%ck

OpenStudy (fibonaccichick666):

@eliassaab Can you elaborate on that last bit?

OpenStudy (anonymous):

It is \( 2^{- }\), which means we approach 2 from the left.

OpenStudy (nincompoop):

(-2)^2 -4(-2) + 4 = 4+8+4 = not zero

OpenStudy (fibonaccichick666):

no its \((-2)^{-}\)

OpenStudy (fibonaccichick666):

right, but the limit is 0/16 if you just plug and chug

OpenStudy (anonymous):

If it is -2 than the answer is zero. if 2, it is +infinity

OpenStudy (fibonaccichick666):

right, I agree it's zero, I need to prove the book wrong

OpenStudy (nincompoop):

graphically confirmed that 2^- = +infinity -2^- = 0

OpenStudy (nincompoop):

the kink is located at -2

OpenStudy (nincompoop):

but I thought that limit DNE at kinks

OpenStudy (fibonaccichick666):

I know there is a cusp, but I just don't know how to show that without being allowed a graphing calculator

OpenStudy (nincompoop):

@satellite73

OpenStudy (fibonaccichick666):

but anyways it's -2 not 2

OpenStudy (anonymous):

\[ |x^2-x-6| = \left\{\begin{array}{rcc} (x-3)(x+2) & \text{if} & (x-3)(x+2) \geq 0 \\ -(x-3)(x+2)& \text{if} & (x-3)(x+2) <0 \end{array} \right.\]

OpenStudy (anonymous):

since you are going to \(-2\) from below, both \((x-3)\) and \((x+2)\) are negative

OpenStudy (anonymous):

and so \(|x^2-x-6|=(x-3)(x+2)\)

OpenStudy (anonymous):

oops that last part is wrong

OpenStudy (anonymous):

you get \[\frac{(x-3)(x+2)}{(x-2)^2}\]

OpenStudy (anonymous):

wait a sec why can't you just plug in \(-2\)? the denominator will not be zero the numerator will you get 0

OpenStudy (anonymous):

oh i see, book says \(\infty\) either there is a typo in the problem or in the answer you just get zero

OpenStudy (nincompoop):

thank you @satellite73

OpenStudy (anonymous):

lol that's what i get for coming in late question was already answered

OpenStudy (nincompoop):

we just needed another person to look at

OpenStudy (nincompoop):

at the problem

OpenStudy (fibonaccichick666):

Thanks guys, I thought there was a type-o. You have confirmed it! I really appreciate everyone's help!